The following labor standards have been established for a particular product:

Standard labor hours per unit of output 4.4 hours
Standard labor rate $ 17.70 per hour

The following data pertain to operations concerning the product for the last month:

Actual hours worked 6,200 hours
Actual total labor cost $ 110,360
Actual output 1,300 units

Required:

a. What is the labor rate variance for the month?

b.  What is the labor efficiency variance for the month?

(Input all amounts as positive values.)

Labor rate variance ____ _____

Labor efficiency variance ____ _____
Answer #1

Labor rate variance = (SR-AR) * AH

= (17.70 - 110360/6200) * 6200

= 620 UNFAVORABLE

Labor efficiency variance = (SH-AH) * SR

= (1300*4.4 - 6200) * 17.70

= 8496 UNFAVORABLE

COMMENT IF YOU FACE ANY ISSUES